I took the Moku and iPad from QIL over to 40m for PLL Measurement at the 40m. (SURF)
Here I summarize the analysis of thermal conduction through 3 ceramic ball bearings supporting a test wafer in Megastat. I considered a baseplate or housing for the ceramic ball bearings at the temperature of the cold plate, the 3 ceramic ball bearings themselves, and the wafer seated on top. This pathway can be broken down into the following components:
1. Thermal resistance across joint from Al baseplate to ball bearing
2. Thermal resistance of bulk of ball bearing
3. Thermal resistance across joint from ball bearing to wafer
These resistances in series can be summed, and there exist 3 of these pathways in parallel (1 for each bearing) and the net resistance can then be found using the corresponding formula.
Attachment 1 is a plot of the thermal conductance of alumina ceramic (Al2O3), and it was taken from from this paper. The peak value of 99% Al2O3 agrees with what is listed in the ceramic ball bearing spec sheet: 28 W/mK. I've used these tabulated values to obtain component 2 above.
Attachment 2 is a plot of the thermal conductances across various joints, taken from Ekin. I could not find data for the conductance across Al--Al2O3 pressure joints or Al2O3--Si pressure joints, but I have used the value corresponding to a stainless steel--stainless steel pressure joint as a conservative upper bound. (*When I was calculating non-negligible conduction last Friday, I had forgotten to update the joint conductance value from an Al--Al pressure joint (orange box in Attachment 2). This greatly overestimated the joint conduction, and the ceramic joints should be a few orders of magnitude below this.)
The results of this modeling show that the maximum cooling power delivered via conductive contact is below 7e-5 W. Compared to the 6e-3 W of radiative cooling power (~1%), this is negligible. Attachment 3 shows the nominal difference in wafer cooldown - the solid and dashed orange curves are barely distinguishable (T difference of ~0.2K at steady state - RTD uncertainty is +- 0.15K). We should proceed with this design plan for emissivity testing in Megastat.
RTD Calibration Plan Developed on 16-06-2022:
Prepare an ice bath and liquid nitrogen bath to calibrate the RTD.
Ice bath: fill cup (?) with ice. Fill with water until 2 inches below the top of the ice. Let sit for 2 minutes before calibration.
We will use a DMM to measure the resistance across the leads. Right now, we are thinking that we will connect the leads to the resistor using alligator clips or solder them together.
Use linear fit to calibrate the RTD value as a function of resistance using these two reference values.
Repeat for all three RTDs
Q: What is the tolerance of the resistor?
100 ohm +/- .06%. Means that our calculation if done correctly will be between 99.4 → 100.6 ohms in the ice bath. Our RTD is class A, which has tolerance +/- .15 degrees. Pt-100 SHOULD be 100 ohms at 0C – the temp changes linearly.
From the plot, it seems like the slop is ~50 Ohms / 120 K. So a difference of 5 K corresponds to ~2 Ohms.
It could be that your measurement of the resistance is off slightly due to the lead resistances, etc. Are you using a 3-wire or 4-wire probe?
You can get a higher accuracy relative calibration by dunking all the RTDs together in ice water and then in boiling water.
I looked closer at the RTDs, and they are all PT100s with the same resistance-temperature curve [Attachment 1]. So we might need to re-calibrate them.
The chamber is currently warming up, and I plan to cool down again on Monday to a) obtain data for the entire cooldown and b) see if this inner shield behavior is repeated.
I encountered several puzzles in the data from this cooldown (Attachment 1):
1. Due to residual challenges with getting the Megastat SLOW channels recorded to frames (after flood), we've been using a USB to extract data form the CTC100. However, only ~48 hours of cooldown were recorded on the drive - we lost data from the rest of the cooldown. I'm hoping this is due to storage limits on the USB, and if so I will decrease the sampling rate of the CTC100 and/or source a USB with larger storage. Of course this is a temporary solution, and efforts would be better spent on getting data recorded in frames. For now, since every component had reached steady state (except the test mass), most parameters can theoretically still be inferred from the data we obtained. (Note that data logging started a bit after the cryocooler was switched on.)
2. The inner shield appears to have reached a steady-state temperature lower than that of the cold head. This is bizzare, and non-physical - the cold head should be the component supplying cooling power to the rest of the system. Our current Megastat model would not predict such behavior under any choice of parameters. I replaced the inner shield RTD before this cooldown, so I am hoping the new RTD has a different calibration curve than what the CTC100 was expecting. If this is the case, I should be able to use the data to undo the existing calibration and obtain correctly-calibrated temperature for the inner shield. It is worth noting that the new RTD accurately measured room temperature, so if this theory is correct then the calibration error only exists at low temperatures.
On Tuesday 6/7 we met to discuss next steps for getting PD testing up and running in CAML (cryo auxiliary mariner lab). Here is a rough 2-month plan:
Step 0 [1 week]: Get CAML workstation running and confirm connection to EPICS/CDS.
- Organize BNC cables coming in from QIL and connect to the PD testing table.
Step 1 [1 week]: Outline desired measurements and sketch diagram of electronics/components.
- Decide how we want these component organized; i.e. how many boxes, number of inputs/outputs per box.
Step 2 [2 weeks]: Implement desired setup.
Step 3 [2 weeks]: Replicate QE measurements on JPL PDs
- Tweak setup as necessary
Step 4 [2 weeks]: Test MCT detectors
The cryocooler was turned off 6/8 at 11am.
I've modeled the theoretical "best" achievable configuration in Megastat, replacing the flexible strap with a solid copper element, and leaving all but one aperture open in the inner shield [Attachment 1]. In this configuration, the wafer can reach 123K in ~20 hours.
- The difference in time constant between the copper bar end and cold head (purple and blue curves) is due to the thermal resistance across the bar and 2 joints, plus realistic radiative losses from bar.
- The difference in time constant between the cold plate and copper bar end (brown/green and purple) is the thermal resistance across the additional solid copper finger and 2 joints.
- The stead-state temperature difference between the inner shield and the wafer (orange and green) is maintained by heat leaking from 1 aperture in the shield. (We've closed up 4, but best case there will need to be 1 open for the copper bar and leads.
Attachment 2 shows power budgeting of heating and cooling power delivered to various components: 1. copper bar, 2. cold plate, 3. inner shield, 4. Si wafer. The points of intersection of the heating and cooling power curves correspond to the steady-state temperature for each component, which can be verified by Attachment 1. The total heat load on the copper bar and cold plate is just over 70W.
The ideal model shows that there are two areas of potential gains:
Analyzing the most recent cooldown data will hopefully validate the efforts of closing up apertures and painting the entire wafer enclosure with aquadag.
I bought a new photodiode for the West Bridge Labs. It will be housed in the QIL (WB265B) in the central cupboards. There is a QR code on it linking to this page.
[Stephen, Radhika, Hiya]
This entry describes efforts towards the most recent Megastat cooldown with:
1.) the inner shield repainted with aquadag and pieces of foil covering the apertures,
2.) aquadag-painted aluminum foil lining the cold plate,
3.) the original test mass and suspension used in previous cooldowns.
Because of the lead time for ordering/fabricating components for wafer supports, we proceeded to use the previous test mass. This way, the improvements made to the chamber can be quantified by comparing the test mass cooldown to past cooldown data.
On Tuesday 5/31, Hiya and I coated the outside surface of the inner shield with Al foil (previously removed for hole tapping / aquadag repainting - see 2775). We inserted the inner shield into the chamber and cut a circular piece of aquadag-coated Al foil to cover the cold plate [Attachments 1, 2]. A slit was cut in the piece of foil in order to pass it over the copper bar and flexible strap. Then, the foil was lightly rested over the bulk created by the flexible strap and its flange [Attachment 3]. We poked holes in the foil in order to bolt down dog clamps for securing the inner shield; then we tightened these dog clamps all around [Attachment 4].
After completing this, we realized we hadn't applied grease to the bottom lip of the inner shield, for maximizing thermal conductance across the cold plate <--> inner shield joint. We also noticed that the inner shield RTD, previously removed for hole tapping / aquadag repainting, was glitching and read an unphysical temperature value. We wrapped up for the day and made note of these items to address next time.
On Wednesday 6/1, Stephen and I picked up with removing the inner shield and aquadag foil. We applied grease to the bottom lip of the inner shield, in the same manner as previous setups. Stephen noticed that the solder joints of the inner shield RTD (that was glitching) had debonded, and so we re-soldered those joints. We re-inserted the inner shield and aquadag foil and bolted the components down. We placed the suspension frame with the test mass back inside the chamber and connected all RTDs [Attachemnt 5]. Since the heater from previous runs had been damaged, we wrapped the exposed leads with kapton tape and proceeded without a heater. Lastly, a piece of foil was placed under the copper bar opening to cover the cut-out area of the inner shield [Attachment 6].
I closed up the chamber and pumped down on Thursday 6/2 at 2:20pm. The cryocooler was switched on at 2:55pm.
We took newport 1811 PDs, one from CTN lab (suspicious) and one from (I forgot) for their high gain and low dark noise.
The detector diameter is small 0.3mm, but our focusing is sufficient:
The mode field diameter of the PM980 fiber is ~ 6.6um. The beam is collimated by a Thorlabs F240APC-1064 with f = 8.07mm and focused with an f = 30mm lens. It means that the diameter at the focus should be roughly 6.6um*30/8.07 = 0.024mm which is well within the PD active area.
We place the PDs at the focal point of the lens at the BHD readout. The impinging optical power was set to be ~ 0.6mW at each port. In one of the PDs, we measure the DC response with a scope to be ~ 5.5 V/0.6 mW ~ 9e3 V/W. According to the specs, the DC monitor as a response of 1e4 V/A while the responsivity of the PD itself is ~ 0.8 A/W at 1064nm so the overall responsivity is ~ 8e3 V/W.
However, the second PD's DC response was bonkers: we measured it to be ten times less. The AC response might still be OK since it is a different port but we haven't measured it yet.
that's a good start, but we want the cooldown to be fast, so what we would need from the model is for you to change some parameters and find out what kind of possible physical changes we can make to make the cooldown fast. i.e. change some parameters and see what configurations would give a fast cooldown, and then we can discuss which of those is the easiest.
I modeled the cooldown of a 2" Si wafer in Megastat with cooling from a) the existing cryocooler and b) LN2. The only difference between the two models is the temperature that the cold end of the copper bar "sees" - in the cryo-cooler case, I've used cold head temperature data from a previous cooldown; in the LN2 case, I've used 77K. In Attachment 1, the cryo-cooler models are solid lines and the LN2 models are dashed.
It is clear from the plot that the cold head gets ~30K colder than LN2 at 77K. This explains the discrepancy between the inner shield models for the two cases at steady state. While the initial temperature rates of change are greater in the LN2 case, the cold head crosses the 77K line in roughly 5 hours.
Does the true cold-head temperature follow the model? Or is it less good than the model?
Today we further discussed ideas for the enclosure, suspension/support, and heat actuation for wafers in Megastat.
Enclosure: We re-painted the inner surface of the inner shield with aquadag, including the top side of the bottom lip [Attachments 1,2]. The folded sheets of aluminum foil covering the apertures were unbolted and also painted with aquadag [Attachment 3].
Wafer support and heat actuation: Attachment 4 shows sketches and a plan forward (drawn by Stephen). The first round of upgrades includes the components underlined in blue: a baseplate to house the points of contact to the wafer, the ceramic ball bearings which will serve as the points of contact, an aquadag-painted sheet metal insert that will bolt down to the coldplate, and a mount for a Maglite mini flashlight.
We will obtain more wire-wound resistors to serve as a back-up to flashlight heating, if the need arises.
It is clear from the plot that the cold head gets ~30K colder than LN2 at 77K. This explains the discrepancy between the inner shield models for the two cases at steady state. While the initial temperature rates of change are greater in the LN2 case, the cold head crosses the 77K line in roughly 5 hours.
I've modeled the cooldown of a 2" diameter and 4" diameter Si wafer in Attachments 1 and 2, using the current Megastat model and previous cold head temperature data. The model includes heat leaking into the inner shield enclosure from an aperture, which we currently observe in Megastat cooldowns. (Note how the wafer cools down much faster than the current test mass, due to the very tiny volume.)
The analytic equation for radiative heat transfer in a 2-surface enclosure (formed by the inner shield and Si wafer) is:
This is dependent on properties of inner shield / cold plate, and as such the accuracy of wafer emissivity measurements will be limited by our uncertainty on the inner shield and cold plate emissivities.
As the ratio approaches 0, the above equation simplifies to:
. The terms related to the surrounding surface (inner shield) drop out of the equation, and so the smaller the ratio of areas, the less of an impact the inner shield / cold plate emissivities will have on the cooldown. Thus we should seek to minimize the ratio of areas to minimize the uncertainty on eSi.
On the other hand, in this low area ratio limit, the thermal power transfer between the wafer and surrounding inner shield is proportional to the area of the wafer. As the attachments show, the 4" diameter wafer gets colder than the 2". This should be taken into account when determining in what temperature range we would like to fit the wafer emissivity. Larger wafer ---> colder. Do we care about emissivity measurements < 123K? If not, the 2" wafer gets us there.
I think this flashlight would work
2. Lamp source + parabolic reflector
The lamp and reflector would be placed and mounted inside the chamber and directed towards the sample. I found parabolic reflectors in the TCS lab and would need to purchase a suitable light source. Here is an example from my search: https://www.thorlabs.com/newgrouppage9.cfm?objectgroup_id=7541. Its emission is broadband and could probably be modeled as a blackbody, but I would need to look into this further. I don't know how efficient the heat transfer to the sample would be, i.e. how much heating power would actually hit the sample. This would also make the control of heating power much more difficult.
This post will serve as a running list of modifications for Megastat for emissivity estimation (brain dump). It is divided into categories:
Category 1: Inner shield / cold plate
We want to cover the cold plate with Aquadag, either directly or by placing aquadag-coated tiles or Al foil on top of the cold plate. It seems the latter approach is preferable, to avoid directly removing and coating the cold plate. Stephen has prepared aquadag-coated aluminum foil which we will soon assess for this purpose. If for some reason it doesn't seem to be sufficient, we will need to identify/design tiles or chunks of aluminum that we can paint with aquadag and lay on top of the cold plate. While we're coating things with aquadag, there are some spots on the inner shield that could use a touch-up.
Category 2: Suspension of sample
There are several options for suspending a 2" Si wafer or 1" optic.
1. RTD lead suspension
This method would involve wrapping the RTD wires around the sample and somehow hanging or dangling the sample from another surface. The wire would be strain-relieved around another component, like a post, before being varnished to the sample. I will have to play around and try out various configurations to determine if this is feasible / would not strain the RTD contact. This approach would not require additional components in the chamber.
2. Insulating posts
In the case where we cannot achieve RTD lead suspension, we will need to rest the sample on support posts while minimizing conductive heat transfer through said posts. Stephen suggested using ceramic ball bearings bolted down to the cold plate. Further modeling is needed to calculate conduction through these supports. Using a bunch of tiny "pins" was also suggested, but these would need to be similarly modeled, and eventually procured.
Category 3: Heat actuation on sample
There are several options for applying heating power to the suspended sample, each with its own drawbacks.
1. Wire-wound resistor
Binding this directly to the sample will be challenging, especially if the sample is wire-suspended. The resistor will occupy a non-negligible amount of area on the sample, which is not ideal for maximum thermal coupling between the sample and the inner shield. Furthermore, since the emissivity an uncoated Si wafer is quite low (low bulk absorption given thin wafer), the emissivity of the heater body (or the varnish/epoxy) could dominate the coupling and lead to an inaccurate fit for sample emissivity.
3. Laser heating source
A laser could be set up outside Megastat and send its beam through one of the viewports to hit the sample. There would be some reflection from the viewport glass, but the transmission would directly heat the suspended sample without crowding the inside of the chamber. The heating power reaching the sample could more easily be determined and controlled in this method, although a laser would need to be sourced for this purpose. This would also require uncovering an additional viewport, which could contribute more heat leakage into the enclosure.
The results from tapping are in the previous entry, and it seemed to work well. As for the emissivity of the inner surface, it would definitely be better to have it black but I wasn't sure how feasible it would be to coat the coldplate. If this is something we can consider, I won't remove the aquadag from the shield.
I'd recommend drilling a through hole and using a nut, rather than tapping as I suggested earlier. The shield is not thick enough to have many 4-40 threads.
Is it better to have the inner shield inner surface low or high emissivity? \Depending on the effect on the MCMC uncertainties, we may want to make everything black, including the cold plate. i.e. we could mount something like black aquadag tiles on the cold plate.
In order to avoid having to order a new inner shield for emissivity estimation in Megastat (see ), we decided to tap 4 holes around each unused aperture (4) in the inner shield. We used 4-40 screws to secure folded sheets of aluminum foil into each aperture, as seen in Attachments 1-2. (These pieces can be bolted on either the outside or inside surface of the shield, but it is shown here bolted to the outside for clarity.) Attachments 3-4 show a piece of folded foil prepared to conceal the gap created by the slit (cut so that the shield could slide over the copper bar). This can be bolted in every time the inner shield is inserted in the chamber. Lastly, we left open the aperture used for passing though the RTD leads, although eventually we can devise a similar foil block with a narrow opening for the leads to pass through.
The state of the test mass / heater from the last cooldown can be seen in Attachment 5. It looks like the heater "fried" and the internal resistive element came out from the brass casing [Attachment 6]. There were lots of flakes of burnt cigarette paper on the cold plate [Attachment 7] and on the test mass [Attachment 8]. I'll need to investigate this further and make sure the heater still works; otherwise we will need to source a new one.
I took out the suspension frame and cleaned up the coldplate [Attachment 9]. Inevitably some of the Aquadag flakes rubbed into the coldplate, but most of it was removed.
Additionally, the simplest model for determining the emissivity of a suspended sample requires that the entire surface surrounding the sample be of the same emissivity. This way, the surrounding surface can be treated as one component that completely encloses the sample (as opposed to 2 or more surfaces with different emissivities partially enclosing the sample). The cold plate surface is rough aluminum, and so I would like to remove the Aquadag coating from the inner surface of the inner shield to restore it to rough aluminum as well.
I've now used the following model of heat transfer between a suspended Si sample (1), inner shield (2), and aperture heat leak (3) in Megastat:
where the first term is the heat transfer from the inner shield to the sample, and the second term is the heat transfer from an aperture opening to the sample. Note that the second term contains the geometric view factor Fleak, which is dependent on the radii of and separation between the 2 surfaces. (This parameter is also implicit in the first term, but because we approximate the inner shield as nearly completely surrounding the sample, we set this to 1.)
Once again I simulated typical/expected values for the emissivities of silicon and rough aluminum:
, for the sample,
, for the inner shield and outer shield (source of heat leaks).
I further parametrized both Aleak and Fleak in terms of rleak, in the assumption of circular aperture openings. This reduced the number of parameters to consider to 4 (e1, e2, e3, rleak).
As before, I simulated T1 using these emissivities, previous inner shield T data, and the above model. I numerically evaluated dT1/dt from the simulated T1 data.
The resulting evaluations for e1 using the dimensions of a large, cylindrical test mass (like the one we've been cooling in Megastat) can be found in Attachment 1. The uncertainty bounds from e2 and rleak were found by evaluating partial derivatives of e1 with respect to those parameters. (The uncertainty bound from e3 was not even resolvable on the plot, so I excluded it as it is a negligible source of e1 error.) This process is mathematically equivalent to evaluating the 2x2 Fisher matrices formed by e1 and e2, e1 and e3, and e1 and rleak.
The resulting uncertainties can be summarized by:
where I have used reasonable guesses for our uncertainties in these parameters. This tells us that in the high mass/area sample case, the uncertainty in the emissivity of the inner shield is our largest source of potential error in e1 determination, even with heat leaks into the system. This makes sense qualitatively, since the first term in the model dominates and view factor from heat leak apertures is small.
Attachment 2 shows the same analysis but using dimensions of a thin, 2" wafer. The resulting uncertainties can be summarized by:
This tells us that in the low mass/area case, the uncertainty of the size of the heat leak is the largest contributor to error in e1 evaluation. Now, the first term in the model no longer dominates, and heat leaks play a non-negligible roll in the cooldown of a wafer.
I next verified these results by simulating my own form of MC, aka defining a prior distribution on e2 and rleak and observing the posterior on e1. I specified the prior on e2 as a 2-dimensional gaussian (m vs. b) with standard deviations of 1e-4 and 1e-2, respectively. I specified the prior on rleak also as a gaussian with standard deviation 1e-2. The results for both the high mass/area case and low mass/area case can be found in Attachments 3 and 4. These serve to verify the previous error analysis and create groundwork for MCMC analysis.
To design our experimental setup with this analysis in mind, I propose contacting the company/manufacturer of the shields and requesting an inner shield with only 1 or 2 apertures. (1 for the copper linkage to pass through, and maybe 1 for the RTD feed through if they can't be passed in through the first aperture). This can be the inner shield used when we have no use for the optical viewports, AKA for emissivity testing. Since minimizing uncertainty in the emissivity of the inner shield is in our best interest especially in the high mass case, this might also be a good opportunity to specify the roughness/polishing level of the inner shield that we have the best model for. (We could still paint the inner surface of the inner shield with Aquadag, but my only concern with this is not fully being able to quantify Aquadag emissivity vs. temperature.)
Tiles are replaced.
Henry from the Carpentry Shop has started replacing ceiling tiles. They need to be cut to fit each location. There was a lot of set up getting equipment across to Bridge before lunch so not that much progress was made replacing tiles in QIL and TCS Labs. Expect work will speed up as we find our groove.
[Jordan, Jancarlo, Stephen, Radhika]
The Helium leak testing setup from the 40m Bake Lab was transported to WB B265B and setup.
The connection to the cryostat was made to the intrinsic KF25 flex tube of the leak tester, at the location of the KF25 - KF40 adapter.
Leak testing found a strong Helium signature at one location of the bottom end cap o-ring seal. The location has been marked in black marker.
Forensics should continue, first with a
At this time, there is no reason to suspect that
We went to the e-shop to investigate the PD circuits. Completely confused about the behavior of the PDs we decide to gain some sanity by testing a sample AD829 on a breadboard with resistors and capacitors similar to those in the design of the PD circuits shown here. The PD is replaced by a voltage source and a 2kOhm resistor such that 0db gain is expected. We first measure the TF of the opamp with the Moku just with the resistors (attachment 1) then with the compensation capacitors.
We tried powering the opamp with shorting V- to ground like we did in the WOPO lab (for some reason this was how it was connected) and got garbage results (attachment 3).
We then turned to retesting the PD circuits with a proper powering scheme. However, connecting +/-5V and ground from a power supply resulted in the output of the PD circuit being ~ -2V even when the PD is taken out which might suggest that the opamps have really gone bad.
I checked the internet in the TCS/MD Lab and it is running great and the switch is runing properly.
We realized that the PD amp circuit only requires a 5V DC supply so we try that. One of the PD had the right response, although only after cycling the input impedance from 50ohm to 1Mohm which is weird. The other one (which produces the negative signal) was complete bonkers.
We remove the home-built PDs and put 2 Thorlabs PDs (forgot the model) with a bad dark current but a decent response and high saturation current. With these PDs we are limited by the PD noise to about 1.25db od squeezing when 30mW LO is detected on each PD without using electronic amplifiers. Attachment 1 shows the different noise spectra we measured.
We maximize the coupling efficiency before boosting the LO power. For some reason, the coupling between the LO fiber and fiber BS deteriorated but there was no apparent dirt on them upon inspection. We crank up the power and measure PD outputs using the Moku oscilloscope. The PD signals were subtracted digitally, but now we were not able to get the shot noise even after fine-tuning the gains. What went wrong? maybe it's because the PDs have separate power supplies?
Some analysis in this notebook
I've used the following model of heat transfer between a suspended Si sample (1) and the inner shield (2) in Megastat:
, where I have not assumed that A1/A2 << 1.
For this analysis, I simulated temperature data of a sample using models of e1 and e2. I simulated e1 and e2 as linear in T:
, and generated test mass temperature data using these emissivity models and inner shield temperature data from a previous cooldown. My goal was to determine how uncertainty in the emissivity of the inner shield and heat capacity of silicon would propagate to the calculated emissivity of the sample.
I back-calculated the emissivity of the sample using a procedure similar to this paper: https://www.sciencedirect.com/science/article/pii/S0017931019361289?via=ihub. To summarize, I used a Savitzky-Golay (SG) filter in scipy to calculate dTdt from the temperature of the sample, and rearranged the model above to solve for e1.
The uncertainty of e1 can be found by:
I considered Cp_Si and e2 as uncertain parameters of interest, and assumed we do not have significant uncertainty on the geometric parameters such as the areas of the inner shield or the sample, or mass of sample.
The results from this analysis are:
Plots of these uncertainties can be found in Attachments 1 and 2.
Next steps are to add an additional radiative heating term to the model (more realistic given what we see in MS) and repeat this analysis, adding uncertain parameters such as size of heat leak. Transfering this analysis to MCMC is also in progress.
First we turned on the relevant instruments for this experiment after the power shutdown:
- Main laser drivers and doubling cavity controller. We set the current to 2 A as we had it before.
- The waveguide TEC. We tried setting it to 60.99 C (for maximum efficiency) but the temperature ramps up much faster and over shoots the setpoint. So we had to do what we did earlier which was to adiabatically change the setpoint from room temperature and finally set it to something like 63 C so the actual measured temperature stabilizes at ~60.9 C. How do we change the PID parameters on this controller? The settings don't seem to allow for it.
- PD power supply, oscilloscopes, function generator, SR 560s lying nearby
Then we tried to probe further what was going on with the PDs (TL;DR not much made sense or was reproducible):
1. Grabbed 30Hz-3GHz HP spectrum analyzer from the Cryolab. Installed it in the WOPO lab under the optical table. We figured out how to do a zero-span measurement around 10MHz. The SA has only one input so we try to combine the signals with an RF splitter. We test this capability by sourcing the RF splitter with 10MHz 4Vpp sine waves from a function generator and measuring the output with a scope. We measure with the scope 1.44Vpp for each channel. The combined channel was 2.73Vpp. We then realized that we still don't have a way to adjust the gains electronically, so we moved on to trying the RF amplifiers (ZFL500 LN).
We assemble two amps on the two sides of a metal heatsink. We solder their DC inputs such that they are powered with the same wire (Attachment 1). We attach the heatsink to the optical table with an L bracket (Attachment 2).
We powered the amps using a 15V DC power supply and tested them by feeding them with 10MHz 10mVpp sine waves from a function generator. We observe on a scope an amplification by a factor of ~ 22. Which makes a power amplification of ~ 26db consistent with the amplifiers' datasheet.
We couldn't find highpass filters with a cutoff around 1MHz, so we resumed using the DC blocks, we test them by feeding white noise into them with a function generator and observing the resulting spectrum. First, we try the DC blocks with a 50 Ohm resistor in parallel. That happened to just cut the power by half. We ditch the resistor and get almost unity transmission above 20kHz.
Moving on to observing LO shot noise, we open the laser shutter. We find there is only 0.7mW coming out of each port of the fiber BHD BS. We measure the power going into the BS to be 4mW. This means the coupling between the LO fiber and the BS fiber is bad. We inspect the fibers and find a big piece of junk on the BS fiber core. We also find a small particle on the LO fiber side. We cleaned both fibers and after butt coupling them we measure 1.6mW at each port. We raise this power to 2mW per port.
We connect the outputs of the PDs to the amps through the DC blocks. The outputs of the amps were connected to the Moku's inputs. The PDs were responding very badly and their noise was also bad. We bypass the amps to debug what is going on. We connect the PDs to a scope. We see they have 300mV (attachment 3) dark noise which is super bad and that they hardly respond to the light impinging on them (attachment 4). We shall investigate tomorrow.
We shut down the workstations and the FBs by doing sudo shutdown and unplugged them from the wall.
Electronic equipment on the FB rack was shut down and unplugged from the wall.
Diablo's current was ramped down and the control unit was shutdown. Optical table electronic equipment was shutdown and the table's powerstrip was switched off.
Equipment under the optical table was switched off and unplugged.
Below are the outlined steps towards building an MCMC model for estimating the emissivity of a surface inside Megastat, and determining achievable uncertainty bounds:
[Radhika, Chris, Ian, Paco]
There are two CTC100s in QIL, so we will use the following nicknames:
CTCMS: CTC100 recording channels from Megastat
CTC: CTC100 recording channels from IR labs dewar (for PD testing)
After observing that the CTCMS channels were stale / not updating, I tried restarting CTC100.service on qil-nfs. After this, the channel values were blank. Chris then looked into the logs and saw lines such as:
CTCMS C4:CTC-MS_WORKPIECE_TEMP_VAL: No reply from device within 1000 ms
CTCMS was not responding to queries, so I killed CTC100.service so I could manually connect to the device via telnet. The connection was made successfully, but CTCMS was not responding to any commands. I then tried connecting to CTC (the other CTC100 for PD testing) and was able to get responses from it. The commands I sent were:
popup "hello" #creates a pop-up on the CTC100 front panel with text "hello".
*IDN? #returns a string with format: Manufacturer, Model number, serial number, version.
After verifying I was connecting to the right IP and port, Ian and I swapped the ethernet cables plugged into both CTC100 devices and resent the commands. Sure enough, CTC continued responding to queries and CTCMS did not. Since the behavior of the devices didn't change when every connection downstream was swapped, it seems the issue lies somewhere within CTCMS. The configurations might have changed after the device was powered off during flooding, or it was somehow damaged. However, CTCMS is recording data properly and logging to USB, so the issue seems to be elusive.
Paco suggested I update the firmware on CTCMS to ensure it is up-to-date. I emailed Stanford Research Systems with a description of our issue, and requested the updated firmware. I can install the firmware via USB once they get back to me, and/or they might have some ideas of why the device is not responsive. In the meantime, someone more experienced with the CTC100 might want to take a look for signs of damage.
Since the heater seems to not be functional once again (displaying N/A and no power output), I proceeded to turn off the cryocooler so the chamber can warm up over the weekend. Cryocooler was turned off at 11:20am 4/15.
The USB logging from cooldown was extremely buggy, with loads of special characters embedded in the temperature data. After filtering out the corrupted values and plotting, the data still looked jumpy and not reflective of the true temperatures. I can work to further filter/treat the data to be able to use for the model, but we might be better off rerunning this cooldown.
I reformatted the USB (MS-DOS FAT16) and reinserted it into the CTC100 to collect warmup data, but a popup appeared saying the USB was not able to record data. I swapped it out for another USB that I similarly reformatted, and the device accepted it. I remember having issues with the original USB in the past, so I'm hoping the swap fixes the corrupted data issue.
**Update: I extracted a bit of data with the new USB, and it looks pristine! I reinserted it to continue to log warmup data. I'm disposing of the old one.
Restart, without vacuum incursion, of cooldown from QIL/2749. Hopefully we pull data this time, via manual logging while Radhika and Chris figure out how to get the channels uploading again.
- Cryocooler on at ~2:30 pm with Workpiece Temp ~ 250 K.
- Datalogging didn't start till ~5:40 pm because I forgot that manual logging was necessary!
that's good. Can you from these models estimate what the uncertainty will be on the emissivity? i.e. by MCMC or otherwise rather than eyeballing
Quick log describing effort to recover leaky IR Labs dewar.
POC Steve Zoltowski - Stevez@irlabs.com
No fix yet, so I reached out to the vendor for more ideas.
I'll post another log with a summary of what leak(s) we suspect and what the current behavior of the leak is.
Fix Effort 1 - Valve Housing Seal (attachment 1)
The fix IR Labs recommended was to look at the seal of the valve housing.
- There was no sign of any issue with oxidation at any visible location.
- The fluoroelastomer valve seat looks like it has crept (plastically deformed in the shape of the sealing surface underneath) but not dramatically.
- The o-ring looked fine, but I wiped all surfaces and added a bit of Krytox to o-ring and valve seat.
- Photos - https://photos.app.goo.gl/oa4bCxm7xaWRJZDj7
Conclusion: No change to the behavior of the leak.
Fix Effort 2 - Feedthrough Seal
I had not yet explored the seal of the feedthrough to the chamber, except to note that the screws are tight.
- The feedthrough wire leads are plugged in within the chamber, and there is not enough slack on the leads to examine the o-ring. I removed the screws, found I had inadequate access, and replaced the screws.
Conclusion: Cleaning / reseating is deferred.
Fix Effort 3 - Window Seal
I had not yet explored the seal of the window to the chamber, except to note that the screws are tight.
- The window looked ok during removal, and I had no reason to be concerned.
- Removed the o-ring and wiped down o-ring and groove thoroughly with IPA.
- Applied Krytox to chamber-side sealing surface.
- Wiped down chamber sealing surface.
Conclusion: No change to the behavior of the leak
I think this is a nice debugging find. Its not very robust to use the workstations as 24/7 script machines (as we have found out over the years).
Best is to install a conda env on the main framebuilder machine, and run the perpetual scripts there in a tmux session.
Once its all sehup, update the ATF Wiki with a description of haw its done. Workstations crash when users do stuff, so its better if the data gettin script can run as a system service (e.g. systemctl, etc)
On Thursday 3/31 we opened up with a goal to diagnose and fix the heater connection (previously reporting an error). Upon opening, we realized the steel wires suspending the test mass had snapped, and the test mass was sitting on the cold plate [Attachments 1, 2]. The mirror had fallen off one face of the test mass, and the heater had also debonded from the other face [Attachments 3, 4]. Our suspicion was that the wires somehow got sliced by the metal zip tie whose function was to mechanically secure the heater in place. Since it did not serve this function anyway and caused more harm, we decided to ditch the zip tie moving forward.
The vacuum pump was started ~4:15pm on 4/1, followed by the cryocooler at 5pm.
On 4/5, I realized the CTC100 log did not contain any data from the weekend cooldown. I expected it would record the data locally even if the workstations were powered off, but this turned out not to be the case. I turned off the cryocooler at 11:45am, with the heater set to 295K. We will redo the cooldown once the chamber gets close to RT.
During the above investigation, I realized the CTC100 channel values were stale - the values are not being updated and all the channels are showing ~255K. None of the RTDs on the CTC100 front panel were are reporting this temperature, so something is getting in the way of proper telnet connection. The warmup waiting period will give me time to diagnose and debug the issue.
The 13.5" rigid legs would cost $1,891.
we don't really need Pneumatic legs. How much for rigid legs?
The current QIL optical tables are Thor Labs PTH503 (discontinued, replaced by PTH603 - also discontinued) which are 700mm tall and offer a closed pneumatic isolation system (passive isolation). This model is only available at the 700mm height and 600mm, which would only be ~4 inches lower than our current legs.
Newport's equivalent model (SL Series Closed Pneumatic Vibration Isolators without Re-Leveling) comes in a 13.5-inch (343mm) height, which would drop our table height by 14 inches. This would be our ideal table height. The cost for 4 legs would come out to $3,047.
Pictures attached. WS1 and WS2 have been turned back on, since the replacement for the ceiling panels will not arrive for another few weeks according to Facilities.
Facilities will be returning on Monday 4/4 between 8-9 AM to remove all ceiling panels above the workstations in B265B (QIL). Replacement of the panels is not yet scheduled, but in the meantime the open ceiling will be covered and the workstations will still be accessible.
Muddy Waters is not new, but if the facility can fix it we'd take it.
This morning, facilities removed all the porous ceiling panels that had been soaked/damaged by water (in B265B: above WS1 and WS2, see Attachments 1+2; In B265A, see Attachment 3). Specifically in B265A, an enclosure was created (Attachment 4) and a dehumidifier was placed inside. All monitors/equipment underneath the panels were thoroughly covered, and the floors were swept up afterward.
No work was done above the North table in the QIL. I asked about it and facilities said they would look into it, but it wasn't on the schedule for today. A member of facilities also pointed out that the sink in the QIL was running black liquid (Attachment 5). It looks like soil/dirt entered the water pipes? This seemed to also be outside of their scope for today.
Facilities placed a blower and dehumidifier in B265B. I checked the airflow and the air around the tables is comparitively still. The North table is covered and the South table is over pressurized by HEPA filters, so there should be little risk of dust being stirred up.
Flood photo album: https://photos.app.goo.gl/BZAG8DyQzFVTfMNz6 (This link is read-only who has no access to the account)
Some photos of affected areas in B265A and B265B (elog shows some preview photos - click on PDF for full set).
Stephen did a great job cleaning up and drying up. Most equipment is powered off and we're leaving it off for a couple of days to dry completely. We'll want to check the stuff on the red lab cart thoroughly.
When I went into QIL today there was a lot of flooding from water dripping from the ceiling at several places in the lab. Images attached.